OpenStudy (anonymous):

How do you solve 2x + 3y = -7?

OpenStudy (radar):

You can graph it, it is a straight line (linear) with infinite number of soluitions.

OpenStudy (anonymous):

HOW do you graph it? @radar

OpenStudy (radar):

It has a slope (m) of -2/3 and a y intercept of -7/3

OpenStudy (radar):

Make a chart and connect the points. I have already gave you one point (0, -7/3) Another point is when x = 10, y = -9 (10,-9). Now you have two points and can construct the line.

OpenStudy (radar):

Another good point is x=7 y=-7 (7,-7)

OpenStudy (radar):

You will have to draw the line yourself.